The power transmitted between two shafts 3.5 metres apart by a cross belt drive round the two pulleys 600 mm and 300 mm in diameters, is 6 kW. The speed of the larger pulley (driver) is 220 r.p.m. The permissible load on the belt is 25 N/mm width of the belt which is 5 mm thick. The coefficient of friction between the smaller pulley surface and the belt is 0.35. Determine : 1. necessary length of the belt ; 2. width of the belt, and 3. necessary initial tension in the belt.


1
Expert's answer
2021-06-04T04:45:02-0400

"L = \fracπ2 (d_1 − d_2) + 2𝑥 + \frac{(d1−d2)^2}{4𝑥}=\fracπ 2(0.6 − 0.3) +(2 × 3.5) + \frac{(0.6−0.3)^2}{4×3.5}= 8.472m"

"v =\frac{
πd_1N_1}{
60 }=
\frac{π × 0.6 × 220}{
60 }= 6.911m/s"

"α = sin^{−1} (\frac{02.6×−3}{2*5.3}) = 7.387°"

"θ = (180 − 2α) ×
\frac{π}{
180} = (180 − (2 × 7.387)) ×
\frac{π}{
180 }= 3.399 rad"

"\frac{T1}{
T2} = e^{μθ} = e^{0.3×3.399} = 3.286 … … . a"

"P = (T_1 \u2212 T_2) \u00d7 v \u2192 6000 = (T_1 \u2212 T_2) \u00d7 6.911"

"→ T_1 − T_2 =\frac{
6000}{
6.911 }= 868.181N → T_1 = 868.181 + T_2 … … . b"

sub a in b

"\frac{868.181 + T_2}{
T_2} = 3.286 → T_2 = 379.781N"

"T_1 = 868.181 + 379.781 = 1247.926N"

"b = \frac{
T1}{
T′ }= \frac{
1247.926}{
25 }= 49.9mm"

"Tₒ = \frac{
T1 + T2}{
2} =\frac{1247.926 + 379.78}{
2 }= 889.08"